OpenStudy (anonymous):

Okay Find G'(x) they said its simple and to plug in what I have in the interval, but what after ? See question in comment box !!

OpenStudy (anonymous):

\[G'(x) = \int\limits_{0}^{\sqrt{x}} 9t \cos (t ^{10}) dt\]

OpenStudy (anonymous):

I have so far is this